\begin{thebibliography}{99}%
{\small
\bibitem{ara}
M. Araskhan, M. R. Rismanchian, Dimension of the $c$-nilpotent multiplier of Lie algebras. Proc. Indian Acad. Sci. Math. Sci. 126 (2016), no. 3, 353-357.
 \bibitem{ba1} P. Batten, K. Moneyhun, E. Stitzinger, On characterizing nilpotent Lie algebras by their multipliers. Comm. Algebra 24 (1996), no. 14, 4319-4330. 
\bibitem{2}
Yu.A. Bahturin, Identical Relations in Lie Algebras, VNU Science Press, b.v., Utrecht, 1987, Trans-lated from the Russian by Bahturin.

\bibitem{cic} S. Cicalò, W. A. de Graaf, C. Schneider, Six-dimensional nilpotent Lie algebras. Linear Algebra Appl. 436 (2012), no. 1, 163-189. 
\bibitem{el}
J. Burns,  G. Ellis, On the nilpotent multipliers of a group, Math. Z. 226 (1997), no. 3, 405--428.

\bibitem{Gr}W. A. de Graaf, Classification of $6$-dimensional nilpotent Lie algebras over fields of characteristic not 2,
 Algebra 309 (2007) 640-653.
\bibitem{ni} P. Niroomand, F.G. Russo, A note on the Schur multiplier of a nilpotent Lie algebra, Comm. Algebra 39 (2011) 1293-1297.
\bibitem{ni2} P. Niroomand, On dimension of the Schur multiplier of nilpotent Lie algebras. Cent. Eur. J. Math. 9 (2011), no. 1, 57-64.
\bibitem{nin} P. Niroomand, M. Parvizi,  F. G. Russo,  Some criteria for detecting capable Lie algebras. J. Algebra 384 (2013), 36--44
\bibitem{ni20}
P. Niroomand, M. Parvizi, $2$-capability and $2$-nilpotent multiplier of finite dimensional nilpotent Lie algebras. J. Geom. Phys. 121 (2017), 180-185.
\bibitem{ni60}
P. Niroomand, F. Johari, M. Parvizi, Capable Lie algebras with the derived subalgebra of dimension $2$ over an arbitrary field. To appear in Linear Multilinear Algebra, doi.org/10.1080/03081087.2018.1425356. 
\bibitem{sa}
F. Saeedi, H. Arabyani, P. Niroomand, On dimension of Schur multiplier of nilpotent Lie algebras II. Asian-Eur. J. Math. 10 (2017), no. 4, 1750076, 8 pp. 
\bibitem{sal} 
A. Salemkar,  B. Edalatzadeh, M. Araskhan,
Some inequalities for the dimension of the $c$-nilpotent multiplier of Lie algebras. J. Algebra 322 (2009), no. 5, 1575-1585.
\bibitem{sa4}
A. Salemkar, A. Aslizadeh,The nilpotent multipliers of the direct sum of Lie algebras. J. Algebra 495 (2018), 220-232.
\bibitem{shi} 
A.I. Shirshov, On bases of free Lie algebras, Algebra Logika 1 (1) (1962) 14-19.

}
\end{thebibliography}

\documentclass{amsart}
\usepackage{amsfonts}
\usepackage{amsmath,amscd}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{newlfont}
\newcommand{\f}{\frac}

 \newtheorem{thm}{Theorem}[section]
\newtheorem{cor}[thm]{Corollary}
 \newtheorem{lem}[thm]{Lemma}
 \newtheorem{prop}[thm]{Proposition}
\theoremstyle{definition}
 
\theoremstyle{remark}

 \theoremstyle{problem}
 
 \numberwithin{equation}{section}

\begin{document}

\title[Characterizing nilpotent Lie algebras ]
 {Characterizing nilpotent Lie algebras rely on the dimension of their $2$-nilpotent multipliers }

\author[F. Johari]{Farangis Johari}
\author[P. Niroomand]{Peyman Niroomand}

\email{e-mail:farangis.johari@mail.um.ac.ir,farangisjohary@yahoo.com}
\address{Department of Pure Mathematics\\
Ferdowsi University of Mashhad, Mashhad, Iran}

\address{School of Mathematics and Computer Science\\
Damghan University, Damghan, Iran}
\email{niroomand@du.ac.ir, p$\_$niroomand@yahoo.com}

\thanks{\textit{Mathematics Subject Classification 2010.} Primary 17B30; Secondary 17B05, 17B99.}

\keywords{$2$-nilpotent multiplier, isoclinism}

\date{\today}


\begin{abstract}There are some results on nilpotent Lie algebras $ L $ investigate the structure  of $ L $ rely on the study of its $2$-nilpotent multiplier. It is showed that the dimension of  the $2$-nilpotent multiplier of $ L $ is equal to $  \frac{1}{3} n(n-2)(n-1)+3-s_2(L).$ Characterizing the structure of all nilpotent Lie algebras has been obtained for the case $ s_2(L)=0.$ This paper is devoted to the characterization of all nilpotent Lie algebras when $ 0\leq s_2(L)\leq 6.$ Moreover, we show that which of them are $2$-capable.
\end{abstract}

\maketitle
\section{Introduction}
For an $n$-dimensional nilpotent non-abelian Lie algebra $ L,$ it is well-know that the dimension of its Schur multiplier is equal to $ \dfrac{1}{2}(n-1)(n-2)+1-s(L) $ for some $ s(L)\geq 0,$ by a result of \cite[Theorem 3.1]{ni}. There are several papers devoted to investigation of the structure of an $n$-dimensional nilpotent non-abelian Lie algebra $ L$ rely on $s(L).$ The structure of all nilpotent non-abelian Lie algebras $ L $ is obtain when $ s(L)=0,1,2,3$ in \cite{ni,ni2,sa}. These results not only characterize a nilpotent Lie algebra in terms of $ s(L) $ but also they can help to shorten the processes of finding the structure of a nilpotent Lie algebra $ L $ in terms of 
$ t(L)=\dfrac{1}{2}n(n-1)-\dim \mathcal{M}(L) $ (see \cite{ba1,ni2}). 

Let $ L$ be a Lie algebra presented as the quotient of a free Lie algebra $ F$ by an ideal $ R.$ Then the $2$-nilpotent multiplier of $ L, \mathcal{M}^{(2)}(L), $ is isomorphic to $ \dfrac{R\cap F^3}{[R,F,F]}.$ It is a less extent the $c$-nilpotent multiplier $ \mathcal{M}^{(c)}(L)  $ for $ c=2 $ (see \cite{ni20}).\\ The study of the $2$-nilpotent multiplier of Lie algebras can lead to the classification of algebras Lie algebra into the equivalence classes as in the group theory case (see \cite{el}).
It also gives a criterion for detecting  the $2$-capability of Lie algebras.
Recall that a Lie algebra $L$ is said to be $2$-capable provided that $L\cong H/Z_2(H)$ for a Lie algebra $H$.
 
In \cite{ni20}, the second author showed that the dimension of  the $2$-nilpotent multiplier of an $n$-dimensional  non-abelian nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $ m $ is bounded by $  \frac{1}{3} (n-m)
\big{(}(n+2m-2)(n-m-1)+3(m-1)
\big{)}+3.$ Then $\dim \mathcal{M}^{(2)}(L)\leq \frac{1}{3} n(n-2)(n-1)+3  $ and so we have $ \dim \mathcal{M}^{(2)}(L) = \frac{1}{3} n(n-2)(n-1)+3-s_2(L)$ for some $s_2(L)\geq 0.$ The structure of all non-abelian   nilpotent Lie algebras is obtained when $s_2(L)=0$ in \cite{ni20}. The current paper is devoted to obtain the structure of all nilpotent non-abelian Lie algebras $ L $ when $1\leq s_2(L)\leq 6.$ Moreover, we specify which of them are capable.
\section{Preliminaries}
Following to Shirshov in \cite{shi}, for a free Lie algebra $L$ on the set $X=\{x_1,x_2,\ldots \}.$
The  basic commutator on the set $X$ defined inductively.
\begin{itemize}
\item[$(i)$] The generators $x_1,x_2,\ldots, x_n$ are basic commutators of length one and ordered by setting $x_i < x_j$ if $i < j.$
\item[$(ii)$] If all the basic commutators $d_i$ of length less than $t$ have been defined and ordered, then we may define the basic commutators of length $t$ to be all commutators of the form $[d_i, d_j]$ such that the sum of lengths of $d_i$ and $d_j$ is $t,$ $d_i > d_j,$ and if $d_i =[d_s, d_t],$ then $d_j\geq d_t.$ The basic commutators of length $t$ follow those of lengths less than $t.$ The basic commutators of the same length can be ordered in any way, but usually the lexicographical order is used.
\end{itemize}
   The number of all basic commutators on a set $X=\{x_1,x_2,\ldots, x_d\}$ of length $n$ is denoted by $l_d(n)$. Thanks to \cite{2},  we have
   \[l_d(n)=\frac{1}{n}\sum_{m|n}\mu (m)d^{\f{n}{m}},\]
   where $\mu (m)$ is the M\"{o}bius function, defined by $\mu (1) = 1, \mu (k) = 0$ if $k$ is divisible by a square, and
$\mu (p_1 \ldots p_s) = (-1)^s $ if $p_1,\ldots , p_s$ are distinct prime numbers.
Using the the topside statement and looking  \cite[Lemma 1.1]{sal} and \cite{shi}, we have the following.\newline

\begin{thm}\label{13}
Let $ F $ be a free Lie algebra on set $ X,$ then $ F^c/ F^{c+i}$ is an abelian Lie algebra with the basis of all basic commutators on $ X $ of lengths $ c,c+1,\ldots,c+i-1 $ for all $0 \leq i \leq c$. In particular, $ F^c/ F^{c+1}$ is an abelian Lie algebra of dimension $l_d(c),$ where $ F^{c} $ is the $c$-th term of the lower central series of $F$.
\end{thm}
The following theorem improves the result of \cite[Theorem 2.5]{ara} for $ c=2 $ when $ L $ is a non-abelian nilpotent Lie algebra.
\begin{thm}\cite[Theorem 2.14]{ni20}\label{1}
Let $  L$ be an $n$-dimensional nilpotent Lie algebra with the derived subalgebra of dimension $m~ (m \geq 1).$ Then
$\dim \mathcal{M}^{(2)}(L) \leq \frac{1}{3} (n-m)
\big{(}(n+2m-2)(n-m-1)+3(m-1)
\big{)}+3.$ If $ m=1, $ then $ \dim \mathcal{M}^{(2)}(L)=  \frac{1}{3}n(n-1)(n-2)+3 $ if and only if  $ L\cong H(1)\oplus A(n-3). $
\end{thm}
\section{main results}
This section is devoted to obtain new result on the dimension of the  $ 2$-nilpotent multiplier of a non-abelian nilpotent Lie algebra. We are going to obtain the structure of all Lie algebras $ L $ such that $1\leq s_2(L)\leq 6. $\\
We need the following two easy lemmas for the next investigation.
\begin{lem}\label{2} Let $  L$ be an $n$-dimensional nilpotent Lie algebra with the derived subalgebra of dimension $m~ (m \geq 3).$ Then
$\dim \mathcal{M}^{(2)}(L) \leq \frac{1}{3} n
(n-2)(n-1)-2.$
\end{lem}
\begin{proof}
By using Theorem \ref{1} and our assumption, we have
\begin{align*}
&\dim \mathcal{M}^{(2)}(L) \leq \frac{1}{3} (n-m)
\big{(}(n+2m-2)(n-m-1)+3(m-1) \big{)}+3 \leq \\& \frac{1}{3} (n-3)
\big{(}(n+4)(n-4)+3(3-1) \big{)}+3\\& =\frac{1}{3} (n-3)
\big{(}(n+4)(n-4)+6 \big{)}+3= \frac{1}{3} (n^3-3n^2)-\frac{10n}{3}+10+3-2+2\\&=\frac{1}{3} (n^3-3n^2)-5(\frac{2n}{3}-3)-2\leq  \frac{1}{3} n(n-2)(n-1)-2.
\end{align*}
The result is obtained.
\end{proof} 
\begin{lem}\label{3} Let $  L$ be an $n$-dimensional nilpotent Lie algebra with the derived subalgebra of dimension $2.$ Then
$\dim \mathcal{M}^{(2)}(L) \leq \frac{1}{3} n
(n-2)(n-1)+1.$
\end{lem}
\begin{proof}
By invoking Theorem \ref{1}, we have
\begin{align*}
&\dim \mathcal{M}^{(2)}(L) \leq \frac{1}{3} (n-2)
\big{(}(n+2)(n-3)+3 \big{)}+3 =   \frac{1}{3} (n-2)(n^2-n-3)+3\\& = \frac{1}{3} (n^3-3n^2)-(\frac{n}{3}-5)\leq \frac{1}{3} (n^3-3n^2)+\frac{2n}{3}+1= \frac{1}{3} n(n-2)(n-1)+1,
\end{align*}
as required.
\end{proof}
\begin{thm}\cite[Theorem 2.13]{ni20}\label{4}
Let $  L$ be an $n$-dimensional nilpotent Lie algebra and $\dim L^2=1.$ Then $ L\cong H(k)\oplus A(n-2k-1) $ and 
\begin{itemize}
\item[$  (i)$]  $   \mathcal{M}^{(2)}(L)\cong A(\frac{1}{3}n(n-1)(n-2)+3),$ if $ k=1. $
\item[$  (ii)$] $ \mathcal{M}^{(2)}(L)\cong A(\frac{1}{3}n(n-1)(n-2)),$ for all $ k\geq 2 $.
\end{itemize}
\end{thm}
\begin{cor}\label{25}
There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $ m\geq 1 $ such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)+2$ or equally $s_2(L)=1.$
\end{cor}
\begin{proof}
The result follows from Lemmas \ref{2}, \ref{3} and Theorem \ref{4}. 
\end{proof}
By using the notation and terminology of \cite{cic,Gr}, we have
\begin{prop}\label{m2}
The $ 2$-nilpotent multiplier of the Lie algebras 
\[L_{4,3}=\langle x_1, x_2, x_3, x_4\big{|}[x_1, x_2] = x_3, [x_1, x_3] = x_4\rangle,\]
\[L_{5,8}=\langle x_1, x_2, x_3, x_4,x_5\big{|}[x_1, x_2] = x_4, [x_1, x_3] = x_5\rangle\] and
\[L_{5,5}=\langle x_1, x_2, x_3, x_4,x_5\big{|}[x_1, x_2] = x_3, [x_1, x_3] = x_5=[x_2,x_4]\rangle\]
 is abelian of dimension $6,$  $18$ and $17,$ respectively.
\end{prop}
\begin{proof}
Let  $ L\cong L_{4,3}$ and  $F$ be a free Lie algebra on the set
$ \lbrace x_1, x_2\rbrace $ and  $ R=\langle [x_1,x_2,x_2]\rangle+F^4.$
Since $L_{4,3}$ is of class $ 3, $   $ F^4\subseteq R $ and so \[
\mathcal{M}^{(2)}(L_{4,3}) \cong \dfrac{ \langle [x_1,x_2,x_2]\rangle+F^4/F^6}{[\langle [x_1,x_2,x_2]\rangle+F^4,F,F]/F^6}.\]
 Theorem \ref{13} implies  $\dim F^4/F^6=l_2(4)+l_2(5)=3+6=9.  $
 It is easy to see that $ [R,F,F]/F^6=[\langle [x_1, x_2,x_2] \rangle,F,F]+F^6/F^6= \langle [x_1,x_2,x_2,x_1,x_1]+F^6,[x_1,x_2,x_2,x_2,x_1]$  $+F^6,[x_1,x_2,x_2,x_2,x_2]+F^6,[x_1,x_2,x_2,[x_1,x_2]]+F^6\rangle$
  and so
$\dim [R,F,F]/F^6=4.$
 It follows $\dim \mathcal{M}^{(2)}(L_{4,3})=10-4=6.$\\ Now, let $ L\cong L_{5,8}.$ 
Clearly, $ L= \langle x_1, x_2, x_3\big{|}[x_2, x_3]=[x_i,x_j,x_k]=0, 1\leq i,j,k\leq 3\rangle.$ 
Now assume that $F$ is a free Lie algebra on the set
$ \lbrace x_1, x_2, x_3\rbrace $ and  $ R=\langle [x_2, x_3] \rangle+F^3.$
Since $L_{5,8}$ is of class two,   $ F^3\subseteq R $ and so $
\mathcal{M}^{(2)}(L_{5,8}) \cong \dfrac{ F^3/F^5}{[R,F,F]/F^5}.$ \newline
 Theorem \ref{13} implies  $\dim F^3/F^5=l_3(3)+l_3(4)=8+18=26.$
 It is easy to see that $ [R,F,F]/F^5= \langle [x_2,x_3,x_1,x_1]+F^5,[x_2,x_3,x_1,x_2]+F^5,[x_2,x_3,x_1,x_3]+F^5, [x_2,x_3,x_2,x_1]+F^5,  [x_2,x_3,x_2,x_2]+F^5,[x_2,x_3,x_2,x_3]+F^5,[x_2,x_3,x_3,x_1]+F^5,
 [x_2,x_3,x_3,x_2]+F^5,[x_2,x_3,x_3,x_3]+F^5\rangle.$ 
 Use of  jacobi identity on all triples and make some calculations, we obtain that
 \begin{align*}
 &[x_2,x_3,x_1,x_2]=-[x_1,x_2,[x_2,x_3]]+[x_2,x_3,x_2,x_1],\\
 &[x_2,x_3,x_1,x_3]=-[x_1,x_3,[x_2,x_3]]+[x_2,x_3,x_3,x_1],\\
 &[x_2,x_3,x_2,x_3]=-[x_2,x_3,[x_2,x_3]]+[x_2,x_3,x_3,x_2].
   \end{align*} 
   Therefore
 $[R,F,F]/F^5= \langle [x_2,x_3,x_1,x_1]+F^5,[x_2,x_3,x_2,x_1]+F^5,[x_2,x_3,x_2,x_2]+F^5,[x_2,x_3,x_3,x_1]+F^5, [x_2,x_3,x_3,x_2]+F^5,[x_2,x_3,x_3,x_3]+F^5,[x_1,x_2,[x_1,x_3]]+F^5,[x_1,x_2,[x_2,x_3]]+F^5 \rangle$
  and so
$\dim [R,F,F]/F^5$  $=8.$
 It follows $\dim \mathcal{M}^{(2)}(L_{5,8})=26-8=18.$\\
 Let  $ L\cong L_{5,5}$ and  $F$ be a free Lie algebra on the set
$ \lbrace x_1, x_2,x_4\rbrace $ and 
 $ R=\langle [x_1,x_2,x_2],$  $[x_2,x_4,x_1],[x_2,x_4,x_2],[x_2,x_4,x_4],[x_1,x_4,x_1],[x_1,x_4,x_2],[x_2,x_4,x_4],[x_1,x_4]\rangle+F^4$ so $ R/F^6\cong  F^3+\langle [x_1,x_4]\rangle/\langle [x_1,x_2,x_1]\rangle+F^6.$
Since $L_{5,5}$ is of class $ 3, $   $ F^4\subseteq R $ and so \[
\mathcal{M}^{(2)}(L_{5,5}) \cong \dfrac{ F^3/\langle [x_1,x_2,x_1]\rangle+F^6}{[\langle [x_1,x_4]\rangle,F,F]+\langle [x_1,x_2,x_1]\rangle+F^5/\langle [x_1,x_2,x_1]\rangle+F^6}.\]
 Theorem \ref{13} implies  $\dim F^3/F^6=l_3(3)+l_3(4)+l_3(5)=8+18+l_3(5).  $
 It is easy to see that $ [R,F,F]/F^6= \langle [x_1,x_4,x_1,x_1],[x_1,x_4,x_2,x_1],[x_1,x_4,x_2,x_2],$ \newline $[x_1,x_4,x_4,x_1],[x_1,x_4,x_4,x_2],[x_1,x_4,x_4,x_4],[x_1,x_2,[x_1,x_4]],[x_1,x_4,[x_2,x_4]]\rangle+F^5/F^6$
  and so
$\dim [R,F,F]/F^6=8+l_3(5).$
Therefore $\dim \mathcal{M}^{(2)}(L_{5,5})=l_3(3)+l_3(4)+l_3(5)-1-l_3(5)-8=17,$ as required.
\end{proof}
A Lie algebra $L$ is called capable if $L\cong H/Z(H)$ for a Lie algebra $H$. See \cite{nin} for more information on this topic.
\begin{prop}\label{68}
Let $ L $ be a non-capable $n$-dimensional nilpotent Lie algebra of  class $3$ with the derived subalgebra of dimension $2$  and $n\geq 6.$ Then $\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} (n-1)
(n-2)(n-3)+2.$
\end{prop}
\begin{proof}
By \cite[Lemma 4.5, Corollary 4.11 and  Theorem 5.1]{ni60}, $ Z^{*}(L)=L^3\cong A(1) $ and so $ L/L^3\cong H(1)\oplus A(n-4).$ Since 
$ L $ is not $2$-capable, we have $\dim \mathcal{M}^{(2)}(L)=\dim \mathcal{M}^{(2)}(L/L^3)-1=\frac{1}{3} (n-1)
(n-2)(n-3)+2,$ by using Theorem \ref{4} and \cite[Lemma 2.2 and Theorem 3.2]{ni20}. 
\end{proof}
\begin{lem}\label{5} There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $2$ such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)+1$ or equally $ s_2(L)=2. $ 
\end{lem}
\begin{proof}
By contrary, let there be  an $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $2$ such that
$\dim \mathcal{M}^{(2)}(L) =\frac{1}{3} n
(n-2)(n-1)+1.$ Let $ B $ be a one dimensional central ideal of $L$ is contained in $L^2.$ Since $ \dim (L/B)^2=1, $ we have $ \dim \mathcal{M}^{(2)}(L/B)\leq  \frac{1}{3}(n-1)(n-2)(n-3)+3$ by using Theorem \ref{4}. Now \cite[Theorem 2.4]{ara} implies that
\begin{align*}
&\frac{1}{3}(n-2)(n^2-n)+1=\frac{1}{3}n(n-1)(n-2)+1=\dim \mathcal{M}^{(2)}(L)\leq \dim \mathcal{M}^{(2)}(L) +\\&\dim L^3\cap B\leq \dim \mathcal{M}^{(2)}(L/B)+\dim (L/L^2\otimes L/L^2 \otimes B)\leq  \\&\frac{1}{3}(n-1)(n-2)(n-3)+3+(n-2)^2=\frac{1}{3}(n-2)(n^2-n-3)+3.
\end{align*}
If $ n\geq 5, $ then we have a contradiction. Hence, we should have $ n=4, $ and so by looking at all nilpotent Lie algebras $ L $ listed in \cite{Gr} and our assumption, we obtain that $ L\cong L_{4,3}. $  By our assumption, since $ s_2(L)=2,$ $ \dim( \mathcal{M}^{(2)}(L_{4,3}) )=\frac{1}{3} 4 (4-2)(4-1)+1=9.$ On the other hand, by using Proposition \ref{m2}, we have $\dim( \mathcal{M}^{(2)}(L_{4,3}) )=6.$ It  is a contradiction again. Hence, the result follows.
\end{proof}
Let $cl(L)$ be used to denote the nilpotency class of a Lie algebra $L.$
\begin{thm}\label{51} 
There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $2$   such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)$ or equally $ s_2(L)=3.$ 
\end{thm}
\begin{proof}
By contrary, let there be  an $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $2$ such that
$\dim \mathcal{M}^{(2)}(L) =\frac{1}{3} n
(n-2)(n-1).$ Let $ B $ be a one dimensional central ideal of $L$ is contained in $L^2.$ Since $ \dim (L/B)^2=1, $ we have $ \dim \mathcal{M}^{(2)}(L/B)\leq  \frac{1}{3}(n-1)(n-2)(n-3)+3,$ by using Theorem \ref{4}. Now \cite[Theorem 2.4]{ara} implies 
\begin{align*}
&\frac{1}{3}(n-2)(n^2-n)=\frac{1}{3}n(n-1)(n-2)=\dim \mathcal{M}^{(2)}(L)\leq \\&\dim \mathcal{M}^{(2)}(L/B)+\dim (L/L^2\otimes L/L^2 \otimes B)-\dim L^3\cap B\leq  \\&\frac{1}{3}(n-1)(n-2)(n-3)+3+(n-2)^2-\dim L^3\cap B\\&=\frac{1}{3}(n-2)(n^2-n-3)+3-\dim L^3\cap B.
\end{align*}
If $ cl(L)=2, $ then $ L^3=0 $ so $ n\leq 5.$ If $ cl(L)=3, $ then since $B= L^2\cap Z(L)=L^3\cong A(1), $ we have $ n\leq 4.$
Let  $ cl(L)=2. $  Hence, our assumption and  looking at the classification of all nilpotent Lie algebras listed in \cite{Gr} show that $ L\cong L_{5,8}.$
By Proposition \ref{m2}, we have $\dim( \mathcal{M}^{(2)}(L_{5,8} ))=  18.$  It contradicts  our  assumption that
$\dim (\mathcal{M}^{(2)}(L_{5,8}))=20.$ 
 Now, let $ cl(L)=3. $
By a similar way, we have $ L\cong L_{4,3}.$ Using Proposition \ref{m2}, we have $\dim( \mathcal{M}^{(2)}(L_{4,3} ))=6.$ It contradicts our  assumption that
  $\dim (\mathcal{M}^{(2)}(L_{4,3} ))=8.$  Hence, the supposition is false and the statement is true.
\end{proof}
\begin{thm}\label{519f} 
There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $m=2$   such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)-1$ or equally $ s_2(L)=4.$
\end{thm}
\begin{proof}
By contrary, let there be  an $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $2$ such that
$\dim \mathcal{M}^{(2)}(L) =\frac{1}{3} n
(n-2)(n-1)-1$ and $ B $ be a one dimensional central ideal of $L$ is contained in $L^2.$ Since $ \dim (L/B)^2=1, $ we have $ \dim \mathcal{M}^{(2)}(L/B)\leq  \frac{1}{3}(n-1)(n-2)(n-3)+3,$ by using Theorem \ref{4}. Now \cite[Theorem 2.4]{ara} implies 
\begin{align*}
&\frac{1}{3}(n-2)(n^2-n)-1=\frac{1}{3}n(n-1)(n-2)-1=\dim \mathcal{M}^{(2)}(L)\leq \\& \dim \mathcal{M}^{(2)}(L/B)+\dim (L/L^2\otimes L/L^2 \otimes B)-\dim L^3\cap B\leq  \\&\frac{1}{3}(n-1)(n-2)(n-3)+3+(n-2)^2-\dim L^3\cap B\\&=\frac{1}{3}(n-2)(n^2-n-3)+3-\dim L^3\cap B.
\end{align*}
If $ cl(L)=2, $ then $ L^3=0 $ so $ n\leq 6.$ If $ cl(L)=3, $ then since $B= L^2\cap Z(L)=L^3\cong A(1), $  $ n\leq 5.$
Let  $ cl(L)=2. $ Hence, our assumption and  looking at the classification of all nilpotent  Lie algebras listed in  \cite{cic,Gr}, we obtain $ L\cong L_{5,8}, L\cong L_{5,8}\oplus A(1), L\cong L_{6,22}(\epsilon)$ or $ L\cong L_{6,7}^{(2)}(\eta). $
By Proposition \ref{m2} and \cite[Theorem 2.5]{ni20}, we have  $\dim( \mathcal{M}^{(2)}(L_{5,8} ))=  18$  and $\dim( \mathcal{M}^{(2)}(L_{5,8} \oplus A(1)))=30.$ It contradicts our assumption. Now, let $ L\cong L_{6,22}(\epsilon)  $
and
$ B $ be a one dimensional central ideal of $L_{6,22}(\epsilon)$ is contained in $L_{6,22}(\epsilon)^2.$ Since $ \dim (L_{6,22}(\epsilon)/B)^2=1 $ and $ L_{6,22}(\epsilon)/B\cong H(2), $ we have $ \dim \mathcal{M}^{(2)}(H(2))=  20,$ by using Theorem \ref{4}. Now \cite[Theorem 2.4]{ara} implies 
$\dim \mathcal{M}^{(2)}(L_{6,22}(\epsilon))\leq \dim \mathcal{M}^{(2)}(H(2))+\dim (H(2)/H(2)^2\otimes H(2)/H(2)^2 \otimes B)= 20+16=36.
$ Similarly, we have  $\dim \mathcal{M}^{(2)}(L_{6,7}^{(2)}(\eta))\leq 36. $ They contradict our assumption that $ \dim \mathcal{M}^{(2)}(L_{6,22}(\epsilon))=39= \dim \mathcal{M}^{(2)}(L_{6,7}^{(2)}(\eta)).$
 Now let $ cl(L)=3. $
 Hence, by looking at the classification of all nilpotent Lie algebras listed in  \cite{Gr}, we obtain  $ L\cong L_{4,3}, $ $ L\cong L_{4,3} \oplus A(1) $ or $ L\cong L_{5,5}. $ By Proposition \ref{m2} and \cite[Theorem 2.5]{ni20}, $\dim( \mathcal{M}^{(2)}(L_{4,3} ))=6,~ \dim( \mathcal{M}^{(2)}(L_{5,5} ))=  17$  and $\dim( \mathcal{M}^{(2)}(L_{4,3} \oplus A(1)))=12.$
They contradict our assumption that $ s_2(L)=4.$ Hence the result is obtained.
 \end{proof}
\begin{thm}\label{kl89}
Let $  L$ be an $n$-dimensional nilpotent Lie algebra with the derived subalgebra of dimension $m\geq 1.$ Then
\begin{itemize}
\item[$  (i)$]$ \dim \mathcal{M}^{(2)}(L)=  \frac{1}{3}n(n-2)(n-1)+3 $ or equally $ s_2(L)=0$ if and only if  $ L\cong H(1)\oplus A(n-3). $
\item[$  (ii)$]There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $m\geq 1$ such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)+2$ or equally $ s_2(L)=1.$
\item[$ ( iii)$]There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $m\geq 1$ such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)+1$ or equally $ s_2(L)=2.$
\item[$ ( iv)$]There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $m\geq 2$ such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)$ or equally $ s_2(L)=3.$
\item[$ ( v)$]There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $m\geq 1$ such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)-1$ or equally $ s_2(L)=4.$
\end{itemize}
\end{thm}
\begin{proof}
The result follows  from  Theorem \ref{1}, Lemma \ref{2}, Theorem \ref{4}, Corollary \ref{25}, Lemma \ref{5}, Theorems \ref{51} and \ref{519f}.
\end{proof}
\begin{cor}\label{845}
Let $  L$ be an $n$-dimensional nilpotent Lie algebra with the derived subalgebra of dimension $m\geq 2.$ Then $\dim \mathcal{M}^{(2)}(L)\leq \frac{1}{3} n
(n-2)(n-1)-2.$
\end{cor}
\begin{proof}
The result follows  from Theorem \ref{kl89}.
\end{proof}
\begin{thm}\label{5191} 
There is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $m\geq 3$   such that
$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)-2$ or $\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)-3.$ 
\end{thm}
\begin{proof}
By contrary, let there be  an $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $m\geq 3$ such that
$\dim \mathcal{M}^{(2)}(L) =\frac{1}{3} n
(n-2)(n-1)-2.$ Let $ B $ be a one dimensional central ideal of $L$ is contained in $L^2.$ Since $ \dim (L/B)^2\geq 2, $ we have $ \dim \mathcal{M}^{(2)}(L/B)\leq  \frac{1}{3}(n-1)(n-2)(n-3)-2,$ by using Corollary \ref{845}. Now \cite[Theorem 2.4]{ara} implies 
\begin{align*}
&\frac{1}{3}(n-2)(n^2-n)-2=\frac{1}{3}n(n-1)(n-2)-2=\dim \mathcal{M}^{(2)}(L)\leq \dim \mathcal{M}^{(2)}(L) +\\&\dim L^3\cap B\leq \dim \mathcal{M}^{(2)}(L/B)+\dim (L/L^2\otimes L/L^2 \otimes B)\leq  \\&\frac{1}{3}(n-1)(n-2)(n-3)-2+(n-3)^2,
\end{align*}
and so $ n\leq 3,$ which is a contradiction. By a similar way, we can see that there is no  $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $m\geq 3$   such that
 $\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)-3.$ The result follows.
\end{proof}
\begin{thm}\label{519} 
Let  $ L $ be an  $n$-dimensional nilpotent Lie algebra with the derived subalgebra of dimension $2.$  Then
\begin{itemize}
\item[$(i)$]$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)-2$ or equally $s_2(L)=5$ if and only if $ L\cong L_{5,8}$ or $ L\cong L_{4,3}.$ 
\item[$(ii)$]$\dim \mathcal{M}^{(2)}(L)=\frac{1}{3} n
(n-2)(n-1)-3$ or equally $s_2(L)=6$ if and only if  $ L\cong L_{5,5}.$ 
\end{itemize}
\end{thm}
 \begin{proof}
\begin{itemize}
\item[$(i)$]
Let there be  an $n$-dimensional nilpotent Lie algebra $ L $ with the derived subalgebra of dimension $2$ such that
$\dim \mathcal{M}^{(2)}(L) =\frac{1}{3} n
(n-2)(n-1)-2$ and $ B $ be a one dimensional central ideal of $L$ in contained $L^2.$ Since $ \dim (L/B)^2=1, $ we have $ \dim \mathcal{M}^{(2)}(L/B)\leq  \frac{1}{3}(n-1)(n-2)(n-3)+3$ by using Theorem \ref{4}. Now \cite[Theorem 2.4]{ara} implies 
\begin{align*}
&\frac{1}{3}(n-2)(n^2-n)-2=\frac{1}{3}n(n-1)(n-2)-2=\dim \mathcal{M}^{(2)}(L)\leq \\& \dim \mathcal{M}^{(2)}(L/B)+\dim (L/L^2\otimes L/L^2 \otimes B)-\dim L^3\cap B\leq  \\&\frac{1}{3}(n-1)(n-2)(n-3)+3+(n-2)^2-\dim L^3\cap B\\&=\frac{1}{3}(n-2)(n^2-n-3)+3-\dim L^3\cap B.
\end{align*}
If $ cl(L)=2, $ then $ L^3=0 $ so $ n\leq 7.$ If $ cl(L)=3, $ then since $B= L^2\cap Z(L)=L^3\cong A(1), $  $ n\leq 6.$
Let  $ cl(L)=2. $  Hence, by looking at the classification of all nilpotent Lie algebras listed in \cite{cic,Gr,ni60}, we obtain  $ L\cong L_{5,8}, L\cong L_{5,8}\oplus A(1), L\cong L_{5,8}\oplus A(2),L\cong L_{6,22}(\epsilon),L\cong L_{6,22}(\epsilon)\oplus A(1),$  $ L\cong L_{6,7}^{(2)}(\eta),$ $L\cong L_{6,7}^{(2)}(\eta)\oplus A(1),~L\cong L_1$ or $ L\cong L_2.$
By Proposition \ref{m2} and \cite[Theorem 2.5]{ni20}, $\dim( \mathcal{M}^{(2)}(L_{5,8} ))=  18$  and so  $\dim( \mathcal{M}^{(2)}(L_{5,8} \oplus A(1)))=30$ and $\dim( \mathcal{M}^{(2)}(L_{5,8} \oplus A(2)))=50.$ It contradicts our assumption  that $ s_2(L)=5.$ Now, let $L\cong L_{6,22}(\epsilon)$ and  
$ B $ be a one dimensional central ideal of $L_{6,22}(\epsilon)$ is contained in $L_{6,22}(\epsilon)^2.$ Since $ \dim (L_{6,22}(\epsilon)/B)^2=1 $ and $ L_{6,22}(\epsilon)/B\cong H(2), $ we have $ \dim \mathcal{M}^{(2)}(H(2))=  20,$ by using Theorem \ref{4}. Now \cite[Theorem 2.4]{ara} implies 
$\dim \mathcal{M}^{(2)}(L_{6,22}(\epsilon))\leq \dim \mathcal{M}^{(2)}(H(2))+\dim (H(2)/H(2)^2\otimes H(2)/H(2)^2 \otimes B)= 20+16=36
$ and hence 
$\dim \mathcal{M}^{(2)}(L_{6,22}(\epsilon)\oplus A(1)) \leq 66. $ Similarly, we have  
$\dim \mathcal{M}^{(2)}(L_{6,7}^{(2)}(\eta))\leq 36 $ and hence 
$\dim \mathcal{M}^{(2)}(L_{6,7}^{(2)}(\eta)\oplus A(1)) \leq 66.$ They cannot happen because of our assumption that $ s_2(L)=5$.
Also, if  $ L\cong L_1 $ or $ L\cong L_2, $ then let
$ B $ be a one dimensional central ideal of $L$ is contained in $L^2.$ Since $ \dim (L/B)^2=1 $ and $ L/B\cong H(2)\oplus A(1), $ we have $ \dim \mathcal{M}^{(2)}(H(2)\oplus A(1))=  40,$ by using Theorem \ref{4}. Now \cite[Theorem 2.4]{ara} implies 
$\dim \mathcal{M}^{(2)}(L)\leq \dim \mathcal{M}^{(2)}(H(2)\oplus A(1))+\dim (L/L^2\otimes L/L^2 \otimes B)= 40+25=65,$
which  contradicts  our assumption that $ s_2(L)=5.$ Hence we should have $ L\cong L_{5,8}.$
In the case that $ cl(L)=3. $  Hence, by looking  the classification of all nilpotent Lie algebras of dimension $ 4 $ listed in \cite{Gr}, we obtain  $ L\cong L_{4,3}.$ Proposition \ref{m2} implies $\dim( \mathcal{M}^{(2)}(L_{4,3} ))=6$ so $ S_2(L_{4,3})=5. $ By a similar way, there is no a Lie algebra such that $s_2(L)=5$ when $ \dim L\geq 5.$
 \item[$(ii)$]By a similar technique is used in  the proof of part $(i)$,
 we conclude that  $L\cong L_{5,5}.$ The converse holds by Proposition \ref{m2}.
\end{itemize}
 \end{proof}
 \begin{thm}\label{man}
 Let  $ L $ be an  $n$-dimensional nilpotent Lie algebra with the derived subalgebra of dimension $m\geq 1.$ Then
 \begin{itemize}
 \item[$(a)$] $s_2(L)=0  $ if and only if $ L\cong H(1)\oplus A(n-3).$ 
 \item[$(b)$] There is no  $n$-dimensional nilpotent Lie algebra $ L $  such that $ s_2(L)=1,2,4.$ 
 \item[$(c)$] $s_2(L)=3  $ if and only if $ L\cong H(k)\oplus A(n-2k-1) $ for all $ k\geq 2.$
  \item[$(d)$] $s_2(L)=5  $ if and only if $ L\cong L_{4,3} $ or $ L\cong L_{5,8}$
  \item[$(e)$] $s_2(L)=6  $ if and only if   $ L\cong L_{5,5}.$
 \end{itemize}
 \end{thm}
 \begin{proof}
 The result is obtained by using Theorem \ref{4}, Corollary \ref{845},  Theorems \ref{5191} and \ref{519}.
 \end{proof}
Recall from \cite{ni20}, a Lie algebra $L$ is said to be $2$-capable if $L\cong H/Z_2(H)$ for a Lie algebra $H$.
In the following corollary, we speciy which ones of Lie algebras with $0\leq s_2(L)\leq 6 $ are capable.
 \begin{cor}
 Let  $ L $ be an  $n$-dimensional nilpotent Lie algebra with the derived subalgebra of dimension $m\geq 1$ such that $0\leq s_2(L)\leq 5.  $ Then $ L $ is $2$-capable if and only if $ L\cong H(1)\oplus A(n-3),$ $  L\cong L_{4,3}, $  $ L\cong L_{5,5} $ or $ L\cong L_{5,8}.$
 \end{cor}
 \begin{proof}
 By using Theorem \ref{man},  $L$ is isomorphic to one of the Lie algebras $ H(k)\oplus A(n-3),$ for all $k\geq 1, $ $  L_{4,3}, $ $  L_{5,5} $ or $L_{5,8}.$
By invoking \cite[Theorem 3.3]{ni20}, $  H(1)\oplus A(n-3)$ is   $2$-capable. Let $ L\cong L_{4,3}$ and $ B $ be a one dimensional central ideal of $L$ is contained in $L^2.$ Since $ \dim (L/B)^2=1, $ we have $ \dim \mathcal{M}^{(2)}(L/B)\leq 3,$ by using Theorem \ref{4}. Since $\dim \mathcal{M}^{(2)}(L/B) < \dim \mathcal{M}^{(2)}(L)=5,$ \cite[Theorem 3.2]{ni20} implies $ L_{4,3} $ is $2$-capable. By a similar way,  $ L\cong L_{5,5}$ and $ L\cong L_{5,8}$ are $2$-capable. Hence the result follows.
 \end{proof}

\end{document}